# Interpreting token compositionality in LLMs: A robustness analysis

### v2

@@ 7 Conclusion
Several limitations are acknowledged in our paper. First, the WordNet dataset may not fully represent language diversity across all domains. Second, the employed transformer models are decoder-based only and could be subject to biases from their training data. Third, our findings depend on the Benepar parsing model, which may introduce inaccuracies in linguistic analysis. Additionally, the applicability of our results to other languages has not been tested. Future research should address these issues and consider combining languages and parsing models for further validation.
@@ Appendix A Compositionality and Localisation
In the context of a specific transformer-based LM model implementing an interpretation function of an input s, the question which is central to this work is whether the contiguous composition of tokens is reflected within the structure of the transformer-based LMs and its constituent parts, layers $l_{0}...l_{n}$, multi-head attention, feed-forward layers and residual connections, i.e. whether the representations $\mathbf{h}_{i}^{(k)}$ at each layer $l_{k}$ explicitly encode the composition of contiguous tokens $t_{i},t_{i+1}$, and how the model’s components contribute to this encoding.
@@ B.3 Model Specifications and Fine-tuning Parameters
The results in Table 3 provide a comparative overview of various Transformer models used in this study. We used GPT-2 models (released under the Modified MIT License), Gemma-2B (released under the Gemma Terms of Use), Llama3 models (released under the Meta Llama 3 Community License), and Qwen models (released under Apache License 2.0). All models were used for research purposes, specifically for language modelling and text generation in English, while Qwen models provide additional multilingual support, aligning with their intended usage. The models differ in their number of parameters, layers, heads, and feed-forward (FF) dimensions. The number of parameters ranges from 85M for GPT2-small to 7.8B for LLama3-8B. The activation functions and FF dimensions also highlight variations in the internal processing architecture, influencing the models’ performance across different tasks. In addition to these architectural differences, the models were fine-tuned using a consistent set of hyperparameters. The fine-tuning process spanned over three training epochs with a batch size of 16. The learning rate was set to 5e-5, while a weight decay of 0.01 was applied to prevent overfitting. Training logs were generated every 200 steps, with model checkpoints saved every 1000 steps, but limited to retaining only one checkpoint to manage storage efficiently. The evaluation strategy during fine-tuning was set to evaluate at the end of each epoch, and similarly, the model was saved at the end of each epoch as well.
@@ Appendix C Token Reduction Analysis
Table 5 presents an analysis of activation reduction percentages across different LLMs, particularly for the token-to-words case. In this context, the mean represents the average reduction percentages across samples, while the standard deviation indicates the variability of these reductions. The purpose is to assess whether token reduction across models would highly influence the results of CAP.
The results presented in Tables 7,  8, and 9 indicate that while token reduction percentage influences performance degradation, it is not the sole determining factor. Several key observations support this conclusion, which is discussed below.
@@ Appendix D Evaluating Parsing Accuracy and Addressing the Impact of Benepar Parser Errors
The parser’s performance was evaluated using the following metrics:
  • Precision: Proportion of correctly predicted constituents out of all predicted constituents.
  • Recall: Proportion of correctly predicted constituents out of all ground truth constituents.
  • F1-Score: Harmonic mean of precision and recall, providing an overall performance measure.
  • Accuracy: Percentage of sentences where the predicted constituency structure fully matches the ground truth.
@@ E.1 Baseline Performance
Table 4 summarises the baseline performance of the models used in this paper on the three tasks. The results include the accuracy of each model’s original and fine-tuned versions on the test set described in Table 2. Fine-tuning generally improves performance, particularly in the larger models such as Gemma-2B and Llama3-8B, which show notable increases in accuracy in most tasks, except the IDM task.
E.2 Comprehensive CAP Results for All Models and Tasks
@@
Tables 7, 8, 9, 10, 11, and 12 present the reduction in accuracy when applying word-level and phrasal CAP, respectively, across models and the three tasks: Inverse Dictionary Modelling (IDM), Synonym Prediction (SP), and Hypernym Prediction (HP). The results of phrasal-level CAP for Gemma-2B and Llama3-8B are not reported due to the severe degradation in model performance under these conditions, rendering the outputs effectively unusable.
@@
This $\Delta A$ value is expressed in percentage points. For example, $\Delta A=40$ indicates that the model’s accuracy has decreased by 40 percentage points from its original performance, which could represent a change from $A_{o}=100\%$ to $A_{c}=60\%$, or any other pair of accuracies with a 40 percentage point difference.
The tables report $\Delta A$ for different layer positions (1%, 25%, 75%, and 100%) in both Original and Fine-tuned settings, using three CAP protocols: Max, Mean, and Sum. This representation allows for a direct comparison of CAP’s impact across different models and tasks, independent of their baseline performance levels.
@@
Table 7: Performance drop (in percentage points) for GPT2 (small, medium, large), Gemma-2B, Llama3 (3B, 8B, 3B Instruct), and Qwen2.5 (0.5B, 1.5B, 3B) models after applying word-level CAP for the Inverse Dictionary Modelling (IDM) task. Results are reported for different layer positions (1%, 25%, 75%, and 100%) in both Original and Fine-tuned settings, using three CAP protocols: Max, Mean, and Sum.
@@
Table 8: Performance drop (in percentage points) for GPT2 (small, medium, large), Gemma-2B, Llama3 (3B, 8B, 3B Instruct), and Qwen2.5 (0.5B, 1.5B, 3B) models after applying word-level CAP for the Synonym Prediction (SP) task. Results are reported for different layer positions (1%, 25%, 75%, and 100%) in both Original and Fine-tuned settings, using three CAP protocols: Max, Mean, and Sum.
@@
Table 9: Performance drop (in percentage points) for GPT2 (small, medium, large), Gemma-2B, Llama3 (3B, 8B, 3B Instruct), and Qwen2.5 (0.5B, 1.5B, 3B) models after applying word-level CAP for the Hypernym Prediction (HP) task. Results are reported for different layer positions (1%, 25%, 75%, and 100%) in both Original and Fine-tuned settings, using three CAP protocols: Max, Mean, and Sum.

### v3

@@ 7 Conclusion
Several limitations are acknowledged in our paper. First, the WordNet dataset may not fully represent language diversity across all domains. Second, the employed transformer models are decoder-based only and could be subject to biases from their training data. Third, our findings depend on the Benepar parsing model, which may introduce inaccuracies in linguistic analysis. Additionally, while our tasks provide an indirect signal of meaning preservation, incorporating explicit reconstruction tasks in future work could offer complementary insight into how CAP affects the retention of input-level information. Finally, the applicability of our results to other languages has not been tested. Expanding CAP to multilingual settings and testing with alternative parsers or models trained with different positional encodings would further validate the generality of our findings.
@@ Appendix A Compositionality and Localisation
In the context of a specific transformer-based LM model implementing an interpretation function of an input s, the question which is central to this work is whether the contiguous composition of tokens is reflected within the structure of the transformer-based LMs and its constituent parts, layers $l_{0}...l_{n}$, multi-head attention, feedforward layers and residual connections, i.e. whether the representations $\mathbf{h}_{i}^{(k)}$ at each layer $l_{k}$ explicitly encode the composition of contiguous tokens $t_{i},t_{i+1}$, and how the model’s components contribute to this encoding.
@@ B.3 Model Specifications and Fine-tuning Parameters
Table 3 provides a comparative overview of various Transformer models used in this study. We used GPT2 models (released under the Modified MIT License), Gemma-2B (released under the Gemma Terms of Use), Llama3 models (released under the Meta Llama 3 Community License), and Qwen models (released under Apache License 2.0). The used models were mainly pre-trained on English data, with Qwen and LLama models providing additional multilingual support, which is English, German, French, Italian, Portuguese, Hindi, Spanish, and Thai for LLama, and more than 10 languages, including Chinese, English, French, Spanish, Portuguese, Russian, Arabic, Japanese, Korean, Vietnamese, Thai, and Indonesian for Qwen. All models were used for research purposes, specifically for language modelling and text generation in English, aligning with their intended usage. The models differ in their number of parameters, layers, heads, and feedforward (FF) dimensions. The number of parameters ranges from 85M for GPT2-small to 7.8B for LLama3-8B. The activation functions and FF dimensions also highlight variations in the internal processing architecture, influencing the models’ performance across different tasks. In addition to these architectural differences, the models were fine-tuned using a consistent set of hyperparameters. The fine-tuning process spanned over three training epochs with a batch size of 16. The learning rate was set to 5e-5, while a weight decay of 0.01 was applied to prevent overfitting. Training logs were generated every 200 steps, with model checkpoints saved every 1000 steps, but limited to retaining only one checkpoint to manage storage efficiently. The evaluation strategy during fine-tuning was set to evaluate at the end of each epoch, and similarly, the model was saved at the end of each epoch as well.
B.4 Handling of Sequence Reduction and Positional Encoding in CAP
CAP reduces the number of token-level activations from the original input length $K$ to a shorter grouped sequence length $G$, by merging activations corresponding to word-level or phrase-level constituents. This reduction is applied post-token embedding and affects intermediate activations within the transformer, specifically the outputs of residual blocks or their internal components (e.g., attention or feedforward sublayers). From the point of CAP application onward, the model processes a reduced-length sequence of size $G$. This operation does not alter the model’s input embeddings or positional encodings.
Effect of Positional Encoding Schemes.
The impact of this reduction depends on the positional encoding strategy used by the model: (i) GPT2 models use Sinusoidal positional embeddings, where each position index corresponds to a unique learned embedding. While CAP does not alter these embeddings directly, reducing the sequence length at intermediate layers can introduce misalignment between positional indices and semantic content. This may disrupt downstream attention or feedforward computations that assume consistent positional context; (ii) LLaMA, Qwen, and Gemma models use rotary positional encodings (RoPE), which encode position relationally through rotation in embedding space. These relative encodings are more robust to changes in sequence length, and CAP has a milder impact on positional semantics in these models. Nevertheless, changes in sequence structure may still affect how models integrate cross-token context.
Although CAP does not interfere with the model’s input or positional embedding layer, it alters the spatial structure of activations mid-forward pass. This may influence how transformers aggregate information across positions, especially in models with absolute position encoding. Nevertheless, we did not observe severe performance degradation in those models compared to others. We acknowledge this as a potential contributing factor to the observed degradation under CAP and consider it an important area for future study.
Namely, Embedding-level analysis represents a promising direction for future exploration. Although this work evaluates a wide range of models with differing positional encoding schemes, we acknowledge the need for more targeted analysis of how CAP interacts with these embeddings. In particular, it would be valuable to quantify the impact of CAP under controlled conditions that isolate embedding effects. For instance, experiments using fixed or masked positional encodings, or applying CAP to models trained from scratch with alternative positional schemes, could help disentangle the influence of compositional pooling from that of positional structure.
@@ Appendix C Token Reduction Analysis
Table 5 presents an analysis of activation reduction percentages across different LLMs, particularly for the token-to-words case. In this context, the mean represents the average reduction percentages across samples, while the standard deviation indicates the variability of these reductions. While models within a family (e.g., Qwen) share the same tokeniser and vocabulary, the reduction percentages still vary across tasks (e.g., SP vs. HP) because different tasks involve input definitions or prompts with different average sentence lengths and syntactic complexity, which in turn affect how many groupings are formed under CAP. In other words, although the tokeniser is fixed, the number and size of groupable units (e.g., multi-token words or phrases) are input-dependent. The purpose is to assess whether token reduction across models would highly influence the results of CAP.
@@
The results presented in Tables 9,  10, and 11 indicate that while token reduction percentage influences performance degradation, it is not the sole determining factor. Several key observations support this conclusion, which is discussed below.
@@ Appendix D Evaluating Parsing Accuracy and Addressing the Impact of Benepar Parser Errors
The parser’s performance was evaluated using the following metrics: (i) Precision: Proportion of correctly predicted constituents out of all predicted constituents; (ii) Recall: Proportion of correctly predicted constituents out of all ground truth constituents; (iii) F1-Score: Harmonic mean of precision and recall, providing an overall performance measure; and (iv) Accuracy: Percentage of sentences where the predicted constituency structure fully matches the ground truth.
@@ E.1 Baseline Performance
Table 4 summarises the baseline performance of the models used in this paper on the three tasks. The results include the accuracy of each model’s original and FT versions on the test set described in Table 2. Fine-tuning generally improves performance, particularly in the larger models such as Gemma-2B and Llama3-8B, which show notable increases in accuracy in most tasks, except the IDM task.
Task / Input Prompt Model CAP Layer (Type) Prediction (No CAP) Prediction (W/ CAP) Observation / Interpretation
IDM: lacking embellishment or ornamentation is called a: " Qwen2.5-1.5B Layer 8 (TW) plain ornament Prediction shifts from correct to antonymic, likely due to token merging altering polarity.
IDM: remaining after all deductions is called a: " LLaMA3.1-8B Layer 4 (TW) net gain Subtle financial distinction lost; CAP causes confusion between output and intermediate step.
IDM: make an effort or attempt is called a:" Gemma-2B Layer 1 (TP) try <h1> Invalid token generation suggests breakdown in early compositional encoding.
IDM: a formal series of statements showing that if one thing is true something else necessarily follows from it is called a:" GPT2-L Layer 24 (TP) proof form Loss of logical structure leads to a more abstract or vague concept.
SP: "journal" is a synonym of Qwen2.5-1.5B Layer 18 (TW) diary di Output truncated, likely due to disruption in longer multi-token word embedding.
SP: "get" is a synonym of Qwen2.5-0.5B Layer 16 (TW) catch break Semantic drift under CAP; verb meaning shifts from acquisition to interruption.
HP: "voice" is a type of Gemma1-2B Layer 16 (TW) sound noise Precision reduced; CAP causes substitution with a noisier, less neutral concept.
HP: "guama" is a type of LLaMA3.2-3B Layer 12 (TW) tree street The output reflects a contextual domain shift, likely due to token-level confusion post-CAP.
Table 7: Representative examples of model predictions with and without CAP applied at various layers. Examples highlight semantic degradation and conceptual drift caused by TW-CAP or TP-CAP applied to original models.
Task / Input Prompt Model CAP Layer (Type) Prediction (No CAP) Prediction (W/ CAP) Observation / Interpretation
IDM: prepare for eating by applying heat is called a: " GPT2-S Layer 4 (TW) cook heat CAP leads to a shift from action to cause, indicating surface-level generalisation.
IDM: fail to attend an event or activity is called a: " LLaMA3.2-3B Layer 1 (TW) miss catch CAP appears to invert the meaning, suggesting confusion in early compositional buildup.
IDM: general term for any insect or similar creeping or crawling invertebrate is called a:" Gemma-2B Layer 11 (TP) bug un Invalid token generation suggests breakdown in compositional encoding
IDM: an institution of higher education created to educate and grant degrees often a part of a university is called a:" GPT2-S Layer 1 (TP) college regular CAP reduces specificity, misclassifying to a generic adjective.
SP: "one fourth" is a synonym of Gemma1-2B Layer 10 (TW) fourth half CAP merges related quantities but loses precision, leading to broader but incorrect substitution.
HP: "hotel" is a type of Qwen2.5-3B Layer 16 (TW) building room Shift from category to subcomponent suggests CAP disrupted higher-level abstraction.
HP: "hexagon" is a type of Qwen2.5-3B Layer 16 (TW) polygon plane Hierarchical class (shape) replaced by domain (geometry); abstraction misaligned.
Table 8: Representative examples of model predictions with and without CAP applied at various layers. Each example shows the prompt, model, CAP configuration (layer and type), predictions, and qualitative interpretation. All examples applied to fine-tuned (FT) models.
E.2 Qualitative Analysis of CAP-Induced Prediction Shifts.
Tables 7 and 8 present representative examples of predictions from multiple models across all the tasks, before and after CAP is applied. These examples are drawn from inputs that the model originally predicted correctly, allowing us to isolate the effects of CAP perturbations without confounding them with unrelated model failures. Each example specifies the CAP layer, CAP type (token-to-word or token-to-phrase), and the model involved. Table 7 focuses on predictions made by original (non-fine-tuned) models, while Table 8 includes outputs from fine-tuned variants. Observed shifts include truncation of multi-token terms (e.g., “diary” $\rightarrow$ “di”), polarity inversion (e.g., “plain” $\rightarrow$ “ornament”), loss of abstraction (“polygon” $\rightarrow$ “plane”), and domain misalignment (e.g., “tree” $\rightarrow$ “street”).
These qualitative differences provide interpretability insights that complement the aggregate metrics reported earlier. They reveal how CAP affects not only performance but the nature of model outputs, especially in terms of semantic generalisation, abstraction shifts, and lexical precision. While we do not observe a uniform trend across layers or model families, TP-CAP consistently induces more severe semantic degradation. This suggests that as model capacity increases, internal representations may become more sensitive to disruptions from externally imposed syntactic structures, potentially due, as argued in the main paper, to a stronger reliance on learned token-level dependencies that diverge from higher-level compositional groupings. This analysis highlights the nature of semantic and lexical shifts induced by CAP, reinforcing the need for future task-specific fine-tuning strategies that improve robustness to structured representation pooling.
E.3 Comprehensive CAP Results for All Models and Tasks
@@
Figure 3 and Tables 9- 13, and 14 present the reduction in accuracy when applying word-level and phrasal CAP, respectively, across models and the three tasks: IDM, SP, and HP. The results of phrasal-level CAP for Gemma-2B and Llama3-8B are not reported due to the severe degradation in model performance under these conditions, rendering the outputs effectively unusable.
@@
This $\Delta A$ value is expressed in percentage points. For example, $\Delta A=40$ indicates that the model’s accuracy has decreased by 40 percentage points from its original performance, which could represent a change from $A_{o}=100\%$ to $A_{c}=60\%$, or any other pair of accuracies with a 40 percentage point difference. The tables report $\Delta A$ for different layer positions (1%, 25%, 75%, and 100%) in both Original and Fine-tuned settings, using three CAP protocols: Max, Mean, and Sum. This representation allows for a direct comparison of CAP’s impact across different models and tasks, independent of their baseline performance levels.
@@
Table 9: Performance drop (in percentage points) for GPT2 (small, medium, large), Gemma-2B, Llama3 (3B, 8B), and Qwen2.5 (0.5B, 1.5B, 3B) models after applying word-level CAP for the Inverse Dictionary Modelling (IDM) task. Results are reported for different layer positions (1%, 25%, 75%, and 100%) in both Original and Fine-tuned settings, using three CAP protocols: Max, Mean, and Sum.
@@
Table 10: Performance drop (in percentage points) for GPT2 (small, medium, large), Gemma-2B, Llama3 (3B, 8B), and Qwen2.5 (0.5B, 1.5B, 3B) models after applying word-level CAP for the Synonym Prediction (SP) task. Results are reported for different layer positions (1%, 25%, 75%, and 100%) in both Original and Fine-tuned settings, using three CAP protocols: Max, Mean, and Sum.
@@
Table 11: Performance drop (in percentage points) for GPT2 (small, medium, large), Gemma-2B, Llama3 (3B, 8B), and Qwen2.5 (0.5B, 1.5B, 3B) models after applying word-level CAP for the Hypernym Prediction (HP) task. Results are reported for different layer positions (1%, 25%, 75%, and 100%) in both Original and Fine-tuned settings, using three CAP protocols: Max, Mean, and Sum.